OpenStudy (anonymous):

log18base12 = a and log54base24 = b, then prove that ab + 5(a-b) = 1

OpenStudy (shubhamsrg):

lets convert all to same base, say m => log 18 / log 12 = a= (2log3 +log2)/(2log2 +log3) => log 54/ log 24 =b = (3log3 +log2)/(log3 + 3log2) i dont know,,maybe try now..

OpenStudy (shubhamsrg):

lets assume log 2 =x and log 3=y so (2y + x)/(2x +y) =a (3y +x)/(y+3x) =b a-b = (2y^2 + 7xy +3x^2) - (7xy + 3y^2 + 2x^2) / (something) = 2(y^2 -x^2) -3(y^2 -x^2) / (5xy +y^2 + 6x^2) = (x^2 -y^2)/(5xy +y^2 + 6x^2) tired now :P you may try further..